OpenStudy (anonymous):

Been asking for an answer for over half an hour :( HELP! Is 5x+10 the answer? "Classrooms on one side of the science building are all numbered with consecutive even integers. If the first room on this side of the building is numbered x, write an expression in x for the sum of five classroom numbers in a row. Then simplify this expression."

OpenStudy (campbell_st):

so it they are consecutive even numbers and the 1st room = x 2nd room = x + 2 3rd room = x + 4 4th room = x + 6 5th room = you can work this one out ---------- Total just add the like terms... now many x's and sum the numbers hope this helps

OpenStudy (dls):

Been asking for an answer for over half an hour 9 minutes ago. True story.

OpenStudy (hba):

lol

OpenStudy (anonymous):

Thank you campbell_st . That did help a lot :)